y= -2(x^2+10x+29)
I know you can't factor it from here, and i also know i need to plug it into the "Quadratic Formula" i just don't know what the formula is.. Can anyone help me.

Answer by stanbon(75887)

y= -2(x^2+10x+29)
------------
Google would show you the formula, but here it is:
x=[-b+-sqrt(b^2-4ac)]/(2a)
Fo your problem:
x=[-10+-sqrt(100-4(1)(29)]/(2)
x=[-10+-sqrt(-16)]/2
x=[-5+-2i]
x=-5+2i or x=-5-2i
Cheers,
Stan H.